A cattle train left Miami and traveled toward New York. 14 hours later a diesel train left traveling at 45 km/h in an effort to catch up to the cattle train. After traveling for four hours the diesel train finally caught up. What was the cattle train's average? (The answer is 10km/hr; what is step by step solution)?
Answer by checkley77(12569)

D=RT
D=45*4
D=180 KILOMETERS IS THE DISTANCE THE DIESEL TRAIN VELED.
180=X(14+4)
180=X*18
X=180/18
X=10 KMH WAS THE SPEED OF THE CATTLE TRAIN.